On August 1, 2021, Limbaugh Communications issued $40 million of 8% nonconvertible bonds at 102. The bonds are due on July 31, 2041. Each $1,000 bond was issued with 20 detachable stock warrants, each of which entitled the bondholder to purchase, for $60, one share of Limbaugh Communications’ no par common stock. Interstate Containers purchased 20% of the bond issue. On August 1, 2021, the market value of the common stock was $58 per share and the market value of each warrant was $8.

In February 2032, when Limbaugh’s common stock had a market price of $71 per share and the unamortized discount balance was $1 million, Interstate Containers exercised the warrants it held.

Required:

1. Prepare the journal entries on August 1, 2021, to record (a) the issuance of the bonds by Limbaugh and (b) the investment by Interstate.

2. Prepare the journal entries for both Limbaugh and Interstate in February 2032, to record the exercise of the warrants.
Answer #1

Ans.1

Limbaugh (Issuer)

Cash Received = Face Value * Issued Rate

= $ 40,000,000 * 102% = $ 40,800,000

Equity Stock Warrants = Market Price per warrant * Number of warrant * Number of Bonds

= $ 8 * 20 warrant * ($40,000,000/$1,000) Bonds = $ 6,400,000

Discount on Bonds Payable = Bond Payable + Equity Stock Warrant - Cash Received

= $ 40,000,000 + 6,400,000 - $40,800,000 = $ 5,600,000

Date Account Title and Explanation Debit Credit
Cash $ 40,800,000.00
Discount on Bonds Payable $    5,600,000.00
          Bond Payable $ 40,000,000.00
          Equity Stock Warrant $    6,400,000.00
[To record the issuance of Bond]

Interstate (Investor)

Investment in stock warrants = Equity Stock Warrants * 20%

= $ 6,400,000 * 20% = $ 1,280,000

Investment in bonds = Face Value * 20%

= $ 40,000,000 * 20% = $ 8,000,000

Cash Paid = Face Value * Issued rate * 20%

= $40,000,000 * 102% * 20% = $ 8,160,000

Discount on Bond Investment = Investment in stock warrants + Investment in bonds - Cash Paid

= $ 1,280,000 + $ 8,000,000 - $ 8,160,000 = $ 1,120,000

Date Accounts Title and Explanation Debit Credit
Investment in stock warrants $    1,280,000.00
Investment in bonds $    8,000,000.00
         Discount on Bond Investment $    1,120,000.00
          Cash Paid $    8,160,000.00
[To record the Purchase of Bond]

2.

Limbaugh (Issuer)

Cash Received = Exercise Price per warrant * Number of warrants * Number of Bonds * 20%

[$ 60 * 20 warrants *($40,000,000/$1,000) Bonds * 20% ] = $9,600,000

Equity-stock warrants = Total Equity-stock warrants * 20%

= $ 6,400,000 * 20% = $ 1,280,000

Common Stock = Cash Received - Equity-stock warrants

=$9,600,000 + $ 1,280,000 = $ 10,880,000

Date Accounts Title and Explanation Debit Credit
Cash Received $    9,600,000.00
Equity Stock Warrant $    1,280,000.00
          Common Stock $ 10,880,000.00
[To record the exercise of stock warrant]

Interstate (Investor)

Investment in stock warrants =Equity Stock Warrants * 20%

= $ 6,400,000 * 20% = $ 1,280,000

Cash Paid = Exercise Price per warrant * Number of warrants * Number of Bonds * 20%

= [$ 60 * 20 warrants *($40,000,000/$1,000) Bonds * 20% ] = $9,600,000

Investment in common Stock = Investment in stock warrants + Cash Paid

= = $ 1,280,000 + $9,600,000 = $ 10,880,000

Date Accounts Title and Explanation Debit Credit
Investment in common Stock $ 10,880,000.00
           Investment in stock warrants $    1,280,000.00
           Cash Paid $    9,600,000.00
[To record the exercise of stock warrant]

Image of page 20